Problem 1:

Describe the properties of a binomial distribution.

 

Problem 2:

Consider a binomial problem with number of trials n = 8 and probability of success P = 0.3. Find the following probabilities:

  1. P(x ≤
  2.  5) 
  3. P(x < 5)
  4. P(x ≥
  5.  6) 

 

Problem 3:

40% of student assessments from a large course are marked by lecturer A, and the rest by lecturer B. A group of five students from this course shares a flat. Find the probability:

  1. that all have their work marked by lecturer B.
  2. that at least one has work marked by lecturer A.
  3. that exactly one has work marked by lecturer A.


Please show working out.



1
Expert's answer
2022-03-03T12:36:41-0500



Problem 3.

"P(A)= 0.4"


"P(B)= 1-(P(A)"


"=1-0.4"


"=0.6"

Hence;

1).

"P(B)= 1-(P(A)"


"=1-0.4"


"=0.6"


2).

"=1-P(A)"


"=1-0.4"


"=0.6"


3).

"P(A)= 0.4"